# biblio.bib

### v2

@@ Appendix C Proof of Lemma 1
where all addition on indices is done modulo $d$. As discussed in the main text, Eq. (74) implies a group of non-onsite symmetries $G_{\bf{a},\bf{b}}=\langle U^{\bf{a}},U^{\bf{b}}\rangle$. If we assume the SIMPS is normal, then these symmetries form a faithful representation of $Z_{2}\times Z_{2}$, i.e. $U^{\bf{a}}$ and $U^{\bf{b}}$ are both non-trivial operators and they are distinct from each other, even after modding out by phase factors that are independent of ${\bf{a},\bf{b}}$. To see this, recall that $B_{\bf{a},\bf{b}}^{ij}$ being normal means that there exists an $L$ such that $\mathrm{span}\{B_{\bf{a},\bf{b}}^{s_{1}i_{1}}B_{\bf{a},\bf{b}}^{i_{1}i_{2}}\cdots B_{\bf{a},\bf{b}}^{i_{L_{0}}s_{2}}\}_{i_{1},\dots,i_{L_{0}}}=M_{\chi}$ for all $s_{1},s_{2}$, where $M_{\chi}$ is the space of all $\chi\times\chi$ matrices. In the present case, the product $B_{\bf{a},\bf{b}}^{s_{1}i_{1}}B_{\bf{a},\bf{b}}^{i_{1}i_{2}}\cdots B_{\bf{a},\bf{b}}^{i_{L_{0}}s_{2}}$ is proportional to $X^{\alpha_{s_{1},i_{1},i_{2},\dots,i_{L_{0}},i_{s_{2}}}}Z^{\beta_{s_{1},i_{1},i_{2},\dots,i_{L_{0}},i_{s_{2}}}}$ where,
@@
Now we use the fact that $O$ is a symmetric operator. This implies that $U^{\bf{a}}O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle=O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle$, and similarly for $U^{\bf{b}}$, i.e. the state $O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle$ has neutral symmetry charge. Up to $\mathcal{D}^{\prime}_{\vec{j}}$, Eq. (79) can be interpreted as a symmetry twisted state, defined by the insertion of a virtual symmetry representation into the SIMPS, as discussed in Sec. 3.4.3. The twisted state has neutral charge if and only if $P_{\vec{i},\vec{j}}=I$ for all $\vec{i},\vec{j}$ such that $\mathcal{D}^{\prime}_{\vec{j}}|\vec{i}\rangle\neq 0$ (note that this only works if $U^{\bf{a}}$ and $U^{\bf{b}}$ are distinct non-trivial operators). Therefore, we can remove $P_{\vec{i},\vec{j}}$ from Eq. (79) without changing the state, giving,

### v3

@@ Appendix C Proof of Lemma 1
where all addition on indices is done modulo $d$. As discussed in the main text, Eq. (75) implies a group of non-onsite symmetries $G_{\bf{a},\bf{b}}=\langle U^{\bf{a}},U^{\bf{b}}\rangle$. If we assume the SIMPS is normal, then these symmetries form a faithful representation of $Z_{2}\times Z_{2}$, i.e., $U^{\bf{a}}$ and $U^{\bf{b}}$ are both non-trivial operators and they are distinct from each other, even after modding out by phase factors that are independent of ${\bf{a},\bf{b}}$. To see this, recall that $B_{\bf{a},\bf{b}}^{ij}$ being normal means that there exists an $L$ such that $\mathrm{span}\{B_{\bf{a},\bf{b}}^{s_{1}i_{1}}B_{\bf{a},\bf{b}}^{i_{1}i_{2}}\cdots B_{\bf{a},\bf{b}}^{i_{L_{0}}s_{2}}\}_{i_{1},\dots,i_{L_{0}}}=M_{\chi}$ for all $s_{1},s_{2}$, where $M_{\chi}$ is the space of all $\chi\times\chi$ matrices. In the present case, the product $B_{\bf{a},\bf{b}}^{s_{1}i_{1}}B_{\bf{a},\bf{b}}^{i_{1}i_{2}}\cdots B_{\bf{a},\bf{b}}^{i_{L_{0}}s_{2}}$ is proportional to $X^{\alpha_{s_{1},i_{1},i_{2},\dots,i_{L_{0}},i_{s_{2}}}}Z^{\beta_{s_{1},i_{1},i_{2},\dots,i_{L_{0}},i_{s_{2}}}}$ where,
@@
Now we use the fact that $O$ is a symmetric operator. This implies that $U^{\bf{a}}O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle=O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle$, and similarly for $U^{\bf{b}}$, i.e., the state $O_{\vec{j}}|\psi_{\bf{a},\bf{b}}\rangle$ has neutral symmetry charge. Up to $\mathcal{D}^{\prime}_{\vec{j}}$, Eq. (80) can be interpreted as a symmetry twisted state, defined by the insertion of a virtual symmetry representation into the SIMPS, as discussed in Sec. 3.4.3. The twisted state has neutral charge if and only if $P_{\vec{i},\vec{j}}=I$ for all $\vec{i},\vec{j}$ such that $\mathcal{D}^{\prime}_{\vec{j}}|\vec{i}\rangle\neq 0$ (note that this only works if $U^{\bf{a}}$ and $U^{\bf{b}}$ are distinct non-trivial operators). Therefore, we can remove $P_{\vec{i},\vec{j}}$ from Eq. (80) without changing the state, giving,